• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Iodice

    Log entry time 10:45:36 on November28,2008

    Entry number 251164

    This entry is a followup to: 251139

    Followups:

    keyword=Rich Performances

    RICH now at 2200 V back to normal behavior. Nr. of photons increased w.r.t. reference mainly due to T/H timing adjustment.
    Problems we had during the night were apparently due to the thresholds definition from the analysis of the pedestal runs.
    Thresholds were set ONLY 2 sigmas above pedestal mean.
    Now I used 3 sigma (modified newdecodeRICH.C )
    The threshold scan done last night should be repeated with this receipe


    FIGURE 1